# Unveiling the Capabilities of Large Language Models in Detecting Offensive Language with Annotation Disagreement

### v1

@@ Abstract
LLMs are widely used for offensive language detection due to their advanced capability. However, the challenges posed by human annotation disagreement in real-world datasets remain underexplored. These disagreement samples are difficult to detect due to their ambiguous nature. Additionally, the confidence of LLMs in processing disagreement samples can provide valuable insights into their alignment with human annotators. To address this gap, we systematically evaluate the ability of LLMs to detect offensive language with annotation disagreement. We compare the binary accuracy of multiple LLMs across varying annotation agreement levels and analyze the relationship between LLM confidence and annotation agreement. Furthermore, we investigate the impact of disagreement samples on LLM decision-making during few-shot learning and instruction fine-tuning. Our findings highlight the challenges posed by disagreement samples and offer guidance for improving LLM-based offensive language detection.
Disclaimer: The samples presented by this paper may be considered offensive or vulgar.
@@ 1 Introduction
Offensive language on online platforms refers to socially considered inappropriate speech due to its vulgarity, obscenity, or irreverence Johnson and Lewis (2010), often inflicting psychological harm on individuals and society. As a result, the detection of offensive language has garnered significant attention from researchers in related fields Founta et al. (2018); Davidson et al. (2017).
Recently, researchers have started leveraging large language models (LLMs) to both detect and annotate offensive language, attracted by their extensive background knowledge and impressive reasoning capacity Kumar et al. (2024). Some studies Wang et al. (2023a); Huang et al. (2023) have shown that LLMs’ detection performance is comparable to that of humans. However, a critical gap in the current research is the idealization of offensive language classification into clear binary labels ("offensive" or "non-offensive"). This simplification overlooks the samples with annotation disagreement, which is common in real-world datasets Aroyo et al. (2019); Basile (2020); Uma et al. (2021b).
Compared to samples with unanimous agreement, those with annotation disagreement better reflect the complexity of human judgment in identifying offensive language Klenner et al. (2020); Leonardelli et al. (2021). These samples often feature ambiguous expressions, presenting challenges for both annotators and detector models Sandri et al. (2023). For example, in the low-agreement offensive sample ($O^{0}$) in Table 1, the statement links "Democrats" with "terrorizing old folks". While some annotators interpret it as inciting hostility due to its provocative tone, others view it as merely a strong emotional expression rather than offensive, leading to disagreement. This disagreement highlights the subjectivity and complexity involved in evaluating offensive language.
Furthermore, the decision confidence exhibited by LLMs when processing disagreement samples can offer valuable insights into their alignment with human annotators Weerasooriya et al. (2023); Baumler et al. (2023). An ideal detector should demonstrate higher confidence for samples with unanimous agreement, indicating clearer cases, and lower confidence for samples with low human agreement, which are inherently more ambiguous Uma et al. (2021a); Leonardelli et al. (2023). This alignment can further assess the feasibility of using LLMs to assist in offensive language detection and annotation tasks, revealing whether the model effectively captures the complexities and nuances that human annotators encounter.
In this paper, we explore the capabilities of LLMs in detecting offensive language with annotation disagreement. Specifically, we mainly focus on the following two research questions:
RQ1: How do LLMs perform in detecting offensive language with annotation disagreement?
How do disagreement samples affect LLMs during the learning phase?
For the first question, we evaluate the model in a zero-shot scenario from two perspectives: First, we assess the binary offensive language detection performance of various open-source and closed-source LLMs on samples with varying agreement degrees. Then, we analyze the relationship between annotation agreement and LLM confidence to gain deeper insights into the model’s decision-making process. For the second question, we focus on the impact of samples with varying agreement degrees during the learning phase, specifically examining two mainstream techniques: few-shot learning and instruction fine-tuning. We explore how samples with a single annotation agreement degree and different combinations of agreement categories affect the model’s detection performance. The main contributions of this paper are summarized as follows:
  • We provide a novel perspective on using LLMs for offensive language detection, emphasizing the importance of samples with annotation disagreement in research.
  • We evaluate the detection performance of multiple LLMs on disagreement samples, examining both binary accuracy and the relationship between annotators and LLMs.
@@
  • We explore the impact of disagreement samples on LLM decision-making, with a focus on few-shot learning and instruction fine-tuning techniques.
@@ 2.1 Dataset
Given that disagreements may arise from both the inherent ambiguity of the data itself and low-quality annotations, the chosen benchmark dataset should meet two criteria: high annotation quality and the open availability of unaggregated annotation data. To this end, we use the high-quality offensive language corpus, MD-Agreement Leonardelli et al. (2021), as the dataset. This dataset consists of 10,753 samples collected from Twitter, with samples annotated by five trained annotators. In addition to the majority-voted hard labels, the dataset provides soft labels reflecting the annotation agreement. These soft labels are mainly categorized into three levels based on the degree of annotation agreement: unanimous agreement ($A^{++}$, i.e., agreement among all five annotators), mild agreement ($A^{+}$, i.e., four out of five annotators agreeing on the same label), and weak agreement ($A^{0}$, i.e., three annotators agreeing and two disagreeing). Moreover, offensive comments are categorized as $O^{++}/O^{+}/O^{0}$, while non-offensive comments are represented as $N^{++}/N^{+}/N^{0}$, respectively. Several samples with varying degrees of annotation agreement are provided in Table 1 for reference. The statistics of the MD-agreement dataset are shown in Table 2. The accuracy of its annotations has been validated by Sandri et al. (2023), with further details provided in Appendix A.1. This dataset is the official corpus of SemEval 2023 Task 11 Leonardelli et al. (2023) and has been widely utilized by researchers in the field Deng et al. (2023); Mokhberian et al. (2024).
@@ 2.2 Models
To ensure a comprehensive evaluation, we access a range of LLMs, including both closed-source and open-source models: For closed-source models, we examine several widely used LLMs, including GPT-3.5, GPT-4, GPT-4o, GPT-o1111Since the current GPT-o1 API does not allow adjustment of the temperature coefficient, we only present its binary classification performance at a temperature of 1 in Table 3, which is the default setting. , Claude-3.5, and Gemini-1.5. For open-source models, we utilize three LLMs with different parameter sizes: LLaMa-3 8B and 70B, Qwen-2.5 7B and 72B, and Mixtral 8x7B and 8x22B. More information about the versions of LLMs is provided in Appendix A.4.
3 How Do LLMs Perform in Detecting Offensive Language with Annotation Disagreement?
In this section, we assess the performance of LLMs in a zero-shot scenario, focusing on both the binary accuracy in detecting offensive language and the relationship between annotation agreement and LLM confidence. We utilize all the samples from the MD-Agreement dataset for this evaluation and analyze the performance on subsets with different annotation agreements.
@@ 3.1 Evaluation of Binary Classification Performance
To evaluate the binary classification performance, we set the temperature coefficient of the LLM to 0, ensuring the model selects the category with the highest probability. We use accuracy and F1 score for assessment. The prompt template used for offensive language detection is listed in Appendix A.3 The outputs of the LLM are converted into hard predictions, with 1 indicating offensive and 0 indicating non-offensive. The results of LLMs in binary detection of offensive language are shown in Table 3. Based on the results, we can draw the following conclusions:
(1) Overall, in the zero-shot scenario without specific training, existing LLMs can fairly accurately identify samples with unanimous agreement ($A^{++}$). The average accuracy for closed-source and open-source LLMs is 88.28% and 86.07%, respectively. Furthermore, the detection performance of some outstanding open-source models, such as LLaMa3-70B, is now very close to that of widely used closed-source models. This trend underscores the increasing effectiveness of LLMs in detecting offensive language with clear labels.
(2) As the degree of annotation agreement decreases, the detection performance of LLMs consistently declines. For instance, GPT-4o achieves an F1 score of 85.24% for $A^{++}$ samples, but for mild agreement ($A^{+}$) and weak agreement ($A^{0}$) samples, the F1 scores drop by 10.64% and 28.18%, respectively. Notably, for weak agreement samples, the F1 scores of all LLMs fall below 65%. This finding highlights that samples prone to annotation disagreement pose significant challenges for LLMs in offensive language detection tasks.
(3) Larger model sizes generally improve the detection performance of LLMs, but the improvement is not as noticeable for samples with annotation disagreement. For example, in detecting $A^{++}$ samples, LLaMa3-70B outperforms LLaMa3-8B by 10.64%, but for $A^{0}$ samples, the improvement is only 3.11%. Similarly, for Mixtral and Qwen2.5, models with larger parameters (8$\times$22B and 72B) show nearly the same F1 scores as smaller models (8$\times$7B and 7B) on the $A^{+}$ and $A^{0}$ subsets. This suggests that the ability of the LLMs to recognize disagreement samples, which lie at the ambiguous boundary between offensive and non-offensive language, may not improve substantially with an increase in model size.
@@
(4) Furthermore, we evaluate the accuracy of several LLMs for both offensive and non-offensive language with varying degrees of annotation agreement, as shown in Figure 2. The results indicate that for samples with the same degree of annotation agreement, most models demonstrate lower accuracy in identifying non-offensive language compared to offensive language. In particular, LLMs consistently exhibit the lowest detection accuracy for non-offensive samples with the lowest agreement ($N^{0}$), with an average accuracy of only 45.77%. This phenomenon can be attributed to the models’ tendency to classify these disagreement samples as the offensive language for safety reasons, reflecting an over-sensitivity of LLMs in the task of offensive language detection, a conclusion consistent with the findings of Kumar et al. (2024); Zhang et al. (2024). Furthermore, this behavior prevents LLMs from effectively capturing the ambiguous boundary between offensive and non-offensive language, much like human crowdsource annotators, further highlighting their limited capability in handling disagreement samples.
@@ 3.2 Evaluation of Relationship between Agreement Degree and LLM Confidence
We then evaluate the relationship between LLM confidence and human annotation agreement. The self-consistency method Chen and Mueller (2024); Wang et al. (2023b) is applied to measure the model’s confidence by resampling its outputs with varying temperature coefficients. To ensure a robust evaluation, we perform experiments across five temperature settings: 0, 0.25, 0.5, 0.75, and 1. The final soft predictions are obtained by averaging the hard predictions from the outputs. Mean Squared Error (MSE) is used as the metric to evaluate the alignment degree between LLM confidence and annotation agreement, following Uma et al. (2021a); Leonardelli et al. (2023), where a smaller MSE indicates closer alignment. We also use Spearman’s Rank Correlation Coefficient ($\rho$) to assess the statistical correlation between the two. The detailed introduction of metrics are provided in Appendix A.2. The results are presented in Table 4.
@@
Based on the results, we observe that as annotation agreement decreases, the alignment degree between it and the LLM confidence shows a noticeable decline across multiple models. For example, GPT-4o, which demonstrates superior overall performance, has an MSE of about 0.05 for $A^{++}$ samples with unanimous agreement. However, when detecting $A^{0}$ samples with low agreement, the MSE increases to 0.2. The statistical correlation between the model’s confidence and annotation agreement also weakens accordingly. For $A^{++}$ samples, the correlation is generally high ($\rho>0.7$) for closed-source models, while for $A^{+}$ and $A^{0}$ samples with disagreements, the correlation is medium $(0.3<\rho\leq 0.7$) and low ($\rho<0.3$), respectively. These findings suggest that existing LLMs have not fully aligned with human standards for identifying offensive language, particularly for samples with annotation disagreement, where they struggle to capture the ambiguous nature.
Figure 3: Confusion matrix (raw counts and percentage) between soft predictions of GPT-4o (x-axis) and soft labels (y-axis).
We then analyze the self-consistency of the outputs generated by a single model across multiple samplings, aiming to gain deeper insights into its decision-making mechanism. We utilize Cohen’s Kappa ($\kappa$) as the metric, which is widely employed to assess the consistency of observations in classification tasks. The results of several LLMs’ self-consistency are shown in Figure 2. We note that, although the self-consistency of each model decreases as annotation agreement diminishes, the Kappa values between sampling remain above 0.75 even for $A^{0}$ samples, demonstrating good ($0.80.6\leq\kappa<$ 0.8) or even high ($\kappa\geq$ 0.8) consistency. This suggests that LLMs demonstrate high confidence in their offensive language detection decisions, but this can also lead to overconfidence, causing them to overlook the ambiguity of disagreement samples. This is particularly evident in the $A^{0}$ subset, where the consistency between LLMs and human annotators is notably low.
Furthermore, we construct a confusion matrix to analyze the relationship between the model’s soft predictions and the samples’ soft labels. GPT-4o is chosen as the representative model. The resulting confusion matrix is presented in Figure 3. The experimental results further reveal that even GPT-4, which achieves overall performance closest to the human level, tends to output the same prediction across multiple sampling attempts, regardless of the level of annotation agreement in the samples, exhibiting strong overconfidence in its predictions. This phenomenon underscores that relying on a single LLM lacks the diversity of judgment characteristic of human annotations in offensive language detection. Therefore, relying solely on LLMs to replace humans in detection and annotation remains a significant challenge.
@@
4 How Do Disagreement Samples Affect LLMs During the Learning Phase?
In this section, we examine the influence of samples with varying annotation agreement on the performance of LLMs during the learning phase, focusing on few-shot learning and instruction fine-tuning. We explore the effect of both single-level agreement samples and different agreement combinations on model performance.
4.1 Impact of Disagreement Samples on Few-shot Learning
We evaluate GPT-4o’s binary classification accuracy and the alignment between annotation agreement and the model’s confidence during the few-shot learning. To enhance the effectiveness of learning, we construct positive and negative sample pairs for each agreement degree, randomly selected from the MD-Agreement training set. Following the settings of Leonardelli et al. (2021), we assess the impact of both single-category agreement samples and combinations of samples with varying agreement degrees in the prompt, where the combined pairs use their respective single-category sample sets. For example, the marker $w/\;A^{++}$ refers to introducing sample pairs with unanimous agreement $A^{++}$ in the prompt, including one offensive and one non-offensive sample, while $w/\;A^{+/0}$ combines medium agreement $A^{+}$ and low agreement $A^{0}$ sample pairs, totaling four samples. The results for the test set of MD-Agreement and its three subsets with varying annotation agreement degrees are presented in Table 6. From these results, we draw the following conclusions:
(1) Overall, compared to the zero-shot scenario, few-shot learning can improve the model’s binary classification performance and its alignment with human annotations. Notably, the most explicit accuracy improvement, averaging 3.87%, is observed in the medium agreement subset ($A^{+}$), highlighting that few-shot learning enhances the model’s ability to better understand offensive language. However, on low agreement samples ($A^{0}$), introducing few-shot learning increases the MSE. This is because few-shot learning also boosts LLM confidence, leading to more consistent predictions, which do not align well with the more ambiguous human annotations in low-agreement samples.
(2) In the few-shot learning setup with a single annotation agreement category, using samples with annotation disagreement (i.e., $A^{+}$ and $A^{0}$) results in greater performance improvements across all subsets and evaluation metrics compared to using samples with unanimous agreement ($A^{++}$). This is because disagreement samples often represent the "edge cases” in the annotation process, which helps the model learn the decision boundaries that humans use to distinguish between offensive and non-offensive language. As a result, the model becomes more effective at understanding and detecting this complex linguistic phenomenon.
(3) Incorporating combinations with varying disagreement degrees into prompts can enhance model performance, though the effectiveness of different combinations varies. For GPT-4o, combining unanimous agreement samples with disagreement ones, e.g., $w/\;A^{++/0}$ and $w/\;A^{++/+}$, explicitly outperforms using only disagreement samples $w/\;A^{+/0}$. However, including combinations with multiple agreement degrees $w/\;A^{++/+/0}$ does not lead to further performance improvements; instead, the binary classification accuracy across various subsets is lower than that of $w/\;A^{++/0}$. This is because combining different types of samples increases the complexity of the model’s decision-making process, potentially hindering accurate judgments. Therefore, in few-shot learning, appropriately balancing samples with varying degrees of agreement is crucial for LLMs to effectively understand and detect offensive language.
To ensure the reliability of our conclusions, we replicate the experiment using the open-source LLM Qwen2.5-72B. The results show that the conclusions generally align with those of GPT-4o. Detailed results are provided in Appendix B.3.

### v2

@@ Abstract
Large Language Models (LLMs) have become essential for offensive language detection, yet their ability to handle annotation disagreement remains underexplored. Disagreement samples, which arise from subjective interpretations, pose a unique challenge due to their ambiguous nature. Understanding how LLMs process these cases, particularly their confidence levels, can offer insight into their alignment with human annotators. This study systematically evaluates the performance of multiple LLMs in detecting offensive language at varying levels of annotation agreement. We analyze binary classification accuracy, examine the relationship between model confidence and human disagreement, and explore how disagreement samples influence model decision-making during few-shot learning and instruction fine-tuning. Our findings reveal that LLMs struggle with low-agreement samples, often exhibiting overconfidence in these ambiguous cases. However, utilizing disagreement samples in training improves both detection accuracy and model alignment with human judgment. These insights provide a foundation for enhancing LLM-based offensive language detection in real-world moderation tasks.
Disclaimer: The paper contains content that may be profane, vulgar, or offensive.
@@ 1 Introduction
Motivation.
A fundamental challenge in offensive language detection is annotation disagreement—cases where human annotators provide conflicting labels for the same text. Disagreement arises due to differences in individual perception, cultural context, and linguistic ambiguity, making offensive language detection inherently subjective Aroyo et al. (2019); Basile (2020); Uma et al. (2021b). However, prior research predominantly treats this task as a binary classification problem, assuming consensus among annotators and failing to account for the inherent subjectivity in offensive language perception.
While large language models (LLMs) have been extensively applied to offensive language detection Kumar et al. (2024); Huang et al. (2023), existing studies primarily evaluate their performance on datasets with binary labels, overlooking their ability to handle cases where annotators disagree. This oversimplification limits our understanding of how well LLMs align with human judgment in ambiguous cases. Moreover, models may exhibit overconfidence in cases where human annotators themselves are uncertain, raising concerns about their reliability for real-world moderation.
Research Objectives.
To bridge this gap, we systematically evaluate LLMs’ ability to process disagreement samples, analyzing both classification accuracy and model confidence. Through this, we seek to determine whether LLMs can effectively navigate subjective offensive language judgments and align with human reasoning. An ideal model should express high confidence for unanimously labeled cases and lower confidence for ambiguous samples, reflecting their inherent uncertainty Weerasooriya et al. (2023); Baumler et al. (2023); Uma et al. (2021a); Leonardelli et al. (2023). Our study provides insights into whether LLMs capture these nuances or exhibit overconfidence in disagreement cases, which could undermine their trustworthiness in content moderation.
This paper systematically investigates how LLMs handle annotation disagreement in offensive language detection. Specifically, we address the following research questions: (RQ1) To what extent can LLMs accurately detect offensive language in cases of human annotation disagreement? (RQ2) How do disagreement samples shape LLM learning and influence decision-making?
To answer RQ1, we evaluate multiple open-source and closed-source LLMs in a zero-shot setting, analyzing both classification accuracy and the relationship between annotation agreement and model confidence. For RQ2, we examine the impact of disagreement samples in few-shot learning and instruction fine-tuning, assessing how different agreement levels affect model performance.
Contributions.
We summarize the contributions of this paper as follows:
  • We provide the first systematic evaluation of LLMs’ performance in offensive language detection under annotation disagreement, revealing key insights into model reliability and human-AI alignment.
  • We conduct an extensive empirical study on LLMs’ handling of disagreement cases, examining models’ performance, confidence, and alignment with human judgment.
@@
  • We analyze the impact of training on disagreement samples, demonstrating how few-shot learning and instruction fine-tuning on these samples influence LLM decision-making in offensive language detection.
@@ 2.1 Dataset
Since annotation disagreements can stem from both intrinsic linguistic ambiguity and labeling error, selecting an appropriate benchmark dataset requires meeting two key criteria: (1) high annotation quality to ensure reliability, and (2) open access to unaggregated annotations to facilitate fine-grained analysis. To ensure a robust evaluation, we employ the MD-Agreement dataset Leonardelli et al. (2021), a high-quality corpus for offensive language detection. It contains 10,753 tweets, each labeled by five trained human annotators, ensuring a reliable annotation process.
The dataset provides both hard labels (majority-voted labels) and soft labels, which indicate the level of agreement among annotators. The soft labels are categorized into three levels:
  • Unanimous agreement ($A^{++}$): All five annotators agree on the label.
  • Mild agreement ($A^{+}$): Four out of five annotators agree.
  • Weak agreement ($A^{0}$): Only three annotators agree, while two disagree.
Each sample in the dataset is also classified as either offensive or non-offensive, following the same agreement-level framework:
  • Offensive samples ($O^{++}$, $O^{+}$, $O^{0}$): Instances labeled as offensive, where the agreement level corresponds to the unanimous, mild, or weak agreement, respectively.
  • Non-offensive samples ($N^{++}$, $N^{+}$, $N^{0}$): Instances labeled as non-offensive, with the same agreement-level distinctions.
Thus, the agreement notation (+++++ +, $+$, 00) applies uniformly across both offensive and non-offensive categories, ensuring consistency in the dataset’s annotation schema. To facilitate subsequent research, we convert the soft labels into floating-point numbers in the range [0, 1] by averaging the hard labels from five annotators (offensive as 1, non-offensive as 0) for each sample.
Examples of samples across different agreement levels are provided in Table 1, and the overall dataset statistics are presented in Table 2. The reliability of MD-Agreement’s annotations has been independently validated in prior studies Sandri et al. (2023), confirming that the disagreement samples are caused by their inherent ambiguity, rather than labeling errors. MD-Agreement serves as the official corpus for SemEval 2023 Task 11 Leonardelli et al. (2023) and has been widely utilized by researchers in the field Deng et al. (2023); Mokhberian et al. (2024). Further dataset details are provided in Appendix A.1.
@@ 2.2 Models
To ensure a comprehensive evaluation, we include both closed-source and open-source LLMs, covering a range of architectures and parameter sizes. For closed-source models, we evaluate widely used proprietary LLMs, including GPT-3.5, GPT-4, GPT-4o, GPT-o1, Claude-3.5, and Gemini-1.5. We also evaluate three families of open-source LLMs at different scales: LLaMa-3 (8B, 70B), Qwen-2.5 (7B, 72B), and Mixtral (8x7B, 8x22B). Further details on the model versions are provided in Appendix A.4.
3 RQ1: Evaluating LLMs on Offensive Language with Annotation Disagreement
In this section, we evaluate the ability of LLMs to detect offensive language in a zero-shot setting. We focus on two key aspects: (1) binary classification accuracy, assessing how effectively models distinguish offensive from non-offensive language across varying annotation agreement levels, and (2) model confidence, analyzing whether LLMs exhibit appropriate uncertainty in ambiguous cases. These aspects are essential for determining whether LLMs can reliably perform offensive language detection in real-world scenarios, where human annotators often disagree.
@@ 3.1 Evaluation of Binary Classification Performance
We assess binary classification performance by evaluating LLMs in a zero-shot setting without additional fine-tuning. To ensure deterministic predictions, we set the temperature coefficient of the LLMs to 0, forcing the model to select the most probable category. We use accuracy and F1 score as evaluation metrics to measure classification performance. The prompt template used for offensive language detection is provided in Appendix A.3. LLM outputs are converted into hard predictions, where 1 indicates offensive and 0 indicates non-offensive. We utilize all the samples from MD-Agreement for a comprehensive evaluation. The classification results are presented in Table 3. Based on the results, we observe the following key findings:
(1) LLMs achieve high accuracy for unanimous agreement ($A^{++}$) samples.
In the zero-shot setting, LLMs consistently accurately classify unanimously agreed-upon ($A^{++}$) samples, achieving 88.28% accuracy for closed-source models and 86.07% for open-source models. Notably, LLaMa3-70B now performs comparably to proprietary models. These results suggest that LLMs perform well on clear-cut cases, driven by their background knowledge and reasoning capabilities.
(2) LLM performance declines sharply for ambiguous cases.
As annotation agreement decreases, LLMs struggle to classify offensive language consistently. GPT-4o’s F1 score drops from 85.24% on $A^{++}$ samples to 74.6% on $A^{+}$ and 57.06% on $A^{0}$. Similarly, all models score below 65% on $A^{0}$ samples. This highlights LLMs’ inability to resolve subjective cases in the real world, where human disagreement often stems from cultural, contextual, or linguistic nuances that models fail to capture.
(3) Larger models improve accuracy but do not resolve annotation disagreement.
While larger models generally perform better, their improvement shrinks for ambiguous cases. For example, LLaMa3-70B outperforms LLaMa3-8B by 10.64% on $A^{++}$ samples but only by 3.11% on $A^{0}$. Similarly, Mixtral and Qwen2.5 show no substantial gain in detecting disagreement samples despite increased parameters. Model scaling alone does not resolve ambiguity, suggesting that larger models lack the nuanced human reasoning required to navigate subjective cases. Alternative training strategies, such as human-in-the-loop approaches or fine-tuning on disagreement samples, may be necessary.
@@
(4) LLMs are biased toward classifying uncertain cases as offensive.
We evaluate the accuracy for offensive and non-offensive language across different agreement levels, as shown in Figure 1. We observed that across all agreement levels, LLMs demonstrate higher accuracy in identifying offensive language than non-offensive language. In particular, for low-agreement non-offensive samples ($N^{0}$), accuracy drops to 45.77%, indicating a strong tendency to misclassify ambiguous content as offensive. This over-sensitivity could lead to false positives in automated moderation systems, increasing the risk of justified content removal and restricting legitimate speech.
@@ 3.2 Evaluation of Relationship between Agreement Degree and LLM Confidence
We analyze how well LLM confidence aligns with human annotation agreement, as a well-calibrated model should exhibit high confidence for clear cases and lower confidence for ambiguous cases. If LLMs assign high confidence to disagreement samples, this may indicate overconfidence, limiting their ability to reflect human-like uncertainty. To evaluate this, we apply the self-consistency method Chen and Mueller (2024); Wang et al. (2023b), which resamples model outputs under varying temperature settings to estimate confidence.
To measure confidence, we evaluate models under five temperature settings: 0, 0.25, 0.5, 0.75, and 1. Higher temperatures introduce more randomness in predictions, helping assess the model’s certainty across varying conditions. The final confidence score is computed by averaging the hard predictions across these temperature settings.
We use Mean Squared Error (MSE) to measure the alignment between LLM confidence and annotation agreement, where a smaller MSE indicates closer alignment Uma et al. (2021a); Leonardelli et al. (2023). Additionally, we employ Spearman’s Rank Correlation Coefficient ($\rho$) to assess statistical correlation. The detailed metric definitions are provided in Appendix A.2. The results are presented in Table 4.
@@
(1) As annotation agreement decreases, the alignment between model confidence and human agreement weakens.
As annotation agreement decreases, LLMs become less reliable in assessing their own uncertainty. GPT-4o, which performs best overall, has an MSE of 0.05 for $A^{++}$ samples but sees this error rise to 0.2 for $A^{0}$ samples. Additionally, Spearman’s correlation ($\rho$) between confidence and agreement weakens from above 0.7 for unanimous samples to below 0.3 for disagreement cases. This suggests that LLMs do not effectively recognize uncertainty in ambiguous cases. In real-world moderation, this could lead to overconfident misclassifications, where the model assigns a high confidence score to an incorrect label, making it harder to detect errors and apply human oversight.
Figure 3: Confusion matrix (raw counts and percentage) between confidence scores of GPT-4o (x-axis) and soft labels (y-axis).
(2) LLMs demonstrate high self-consistency but may be overconfident in disagreement cases.
We assess self-consistency using Cohen’s Kappa ($\kappa$), measuring how stable LLM outputs remain across multiple sampling attempts. As shown in Figure 2, self-consistency decreases for lower agreement samples but remains above 0.75 even for $A^{0}$ cases, indicating strong internal agreement. While high self-consistency is desirable for clear-cut cases, it becomes problematic in ambiguous cases, as it suggests that LLMs remain overconfident even when human annotators disagree. This rigidity limits the model’s ability to adjust for nuanced linguistic or contextual differences.
(3) Even high-performing models exhibit overconfidence, limiting their ability to reflect human-like uncertainty.
We construct a confusion matrix of GPT-4o to visually analyze the relationship between the model’s confidence score and the soft labels of samples, as shown in Figure 3. The result reveals that even GPT-4o, the best-performing model, assigns high confidence to its predictions regardless of annotation agreement, indicating a lack of adaptability to disagreement cases. This overconfidence highlights a critical flaw in LLM-based moderation: their inability to reflect the diversity of human judgment. Overconfident models are more likely to make systematic errors in handling subjective content, leading to unreliable moderation outcomes. Instead of relying on LLMs as sole decision-makers, future research should explore ensemble methods, uncertainty-aware training, or human-AI collaboration to mitigate biases and improve disagreement resolution.
We further analyze consistency across different models in Appendix B.2, and reveal low agreement among LLMs on disagreement samples. This highlights the potential of ensemble models to handle these nuanced cases.
@@
4 RQ2: Impact of Disagreement Samples on LLM Learning
In this section, we examine how samples with varying annotation agreements influence LLM performance during the learning phase. We focus on two key learning paradigms: few-shot learning and instruction fine-tuning. Specifically, we explore the impact of both single-category agreement samples and different agreement-level combinations on model performance.